OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0. 9-5x^2=(x+2)^2-4

OpenStudy (anonymous):

but what part of 6x^2-5 would be the c part?

OpenStudy (anonymous):

\[9-5x^2=\]\[(x+2)^2-4\] (x+2)^2=\[x ^{2}+4x+4\] \[9-5x ^{2}=x ^{2}+4x+4-4\] \[6x ^{2}+4x-9\] a=6,b=4,c=-9

OpenStudy (anonymous):

i think this may not be right

OpenStudy (anonymous):

Nick's answer is correct

OpenStudy (anonymous):

thanks joe

OpenStudy (anonymous):

oh yes. it just came through. only saw the first one. yes nick is right

OpenStudy (anonymous):

yes yes yes.........nick's ans is crt.........sry..........

OpenStudy (anonymous):

plz reward a medal if its correct.:)